​(​Bondholders' expected rate of return​) You own a bond that has a par value of ​$1000 and matures in 18 years. It pays an annual coupon rate of 8 percent. The bond currently sells for ​$725. What is the​ bond's expected rate of​ return? The​ bond's expected rate of return is nothing​
Answer #1

use financial calculator function as below

FV=1000 is the face value or par value of the bond

PV=-725 is the present value or current price of the bond

N=18

PMT=1000*8%=80

Click CPT

Click I/Y=11.73% is The​ bond's expected rate of return

the above is answer..